At General Hospital, nurses are given performance evaluations to determine eligibility for merit pay raises. The supervisor rates the nurses on a scale of 1 to 10 (10 being the highest rating) for several activities: promptness, record keeping, appearance, and bedside manner with patients. Then an average is determined by giving a weight of 2 for promptness, 4 for record keeping, 1 for appearance, and 3 for bedside manner with patients. What is the average rating for a nurse with ratings of 8 for promptness, 9 for record keeping, 3 for appearance, and 1 for bedside manner? (Enter your answer to one decimal place.)
Answer #1

GIVEN:

The ratings for a nurse is as follows:

8 for promptness, 9 for record keeping, 3 for appearance, and 1 for bedside manner

Their corresponding weights are given by:

2 for promptness, 4 for record keeping, 1 for appearance, and 3 for bedside manner with patients

TO FIND:

Average rating for a nurse = ?

FORMULA USED:

The formula for weighted average is given by,

Weighted average =sum xw/sum w

where x is a data value and w is the weight assigned to that data value.

CALCULATION:

Rating (x)    Weight (w) Weighted value rw
   8    2    8*2 = 16
   9    4    9*4 = 36
   3    1    3*1 = 3
   1    3    1*3 = 3

We know that

Weighted average =sum xw/sum w

  (16 +36+3+3)/(2+4+1+3)

  58/10

Weighted average =5.8

Thus the average rating for a nurse with ratings of 8 for promptness, 9 for record keeping, 3 for appearance, and 1 for bedside manner is 5.8.